OpenStudy (goformit100):

Let U = {1, 2, 3, 4, 5, 6}, A = {2, 3} and B = {3, 4, 5}. Find A′, B′, A′∩B′, A∪B and hence show that ( A∪B)′= A′∩B′.

OpenStudy (anonymous):

I'm going to copy paste ALL the HW!

OpenStudy (mertsj):

A'=U-A B'=U-B A' intersect B' = elements in both A' and B' AUB= A+B

OpenStudy (goformit100):

Thank you sir

OpenStudy (usukidoll):

I thought intersect is what is unique about A and B.? ? ? ?

OpenStudy (usukidoll):

as in what do they NOT have in common

OpenStudy (usukidoll):

I don't see 2, 4, and 5 in both A and B, I see the number 3 in A and B

OpenStudy (usukidoll):

U and A 2 3 what don't they have in common? 1 4 5 6 U and B 3 4 5 what they don't have in common? 1 2 6

OpenStudy (mertsj):

@UsukiDoll This is America. You have a right to be wrong.